OpenStudy (anonymous):

What is the distance from (6, -8) to the origin? quick medal?

OpenStudy (bibby):

use the distance formula, \(\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) we have (0,0) and (6,-8)

OpenStudy (i_love_my_nieces):

Options

OpenStudy (i_love_my_nieces):

@not.exactly.a.genius

OpenStudy (anonymous):

im not sure how to do it? @bibby @I_love_my_nieces

jimthompson5910 (jim_thompson5910):

Using the formula @bibby wrote out, you will have \[\Large d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] \[\Large d = \sqrt{(6-0)^2+(-8-0)^2}\] \[\Large d = \sqrt{(6)^2+(-8)^2}\] ... ... ... \[\Large d = ???\] Tell me what you get

OpenStudy (i_love_my_nieces):

Is it MC?

OpenStudy (i_love_my_nieces):

I KNOW THE ANSWER IS 70

OpenStudy (bibby):

nope

OpenStudy (i_love_my_nieces):

Yes it is.

OpenStudy (bibby):



OpenStudy (i_love_my_nieces):

Well if I don't have the right answer then what is it?

OpenStudy (anonymous):

10